OpenStudy (anonymous):

y = (1/4)x^2 - (1/2)lnx..over the interval (1, 7e) ...what is the arc length ? need help pleease

OpenStudy (anonymous):

s= integral from a to b sqrt(1+(y prime)^2)dx note y prime is the derivative of your function with respect to x and your integral goes from 1 to 7e

OpenStudy (anonymous):

or do you need to know how to do the derivatives and integrals?

OpenStudy (anonymous):

The arclength is given by: \[\int\limits_a^b \sqrt{1+[f'(x)]^2}dx\] So you have: \[f'(x)=(1/2)x-1/(2x) \rightarrow [f'(x)]^2=((1/2)x-1/(2x))^2=((1/4)x^2+1/(4x^2)-(1/2))\] Then you have: \[\int\limits_1^{7e} \sqrt{1+(1/4)x^2+1/(4x^2)+(1/2)}dx\] As long as my mental squaring is correct.

OpenStudy (anonymous):

i have ((x^2)/4)- ((x^2)/4)+(1/(4x^2))

OpenStudy (anonymous):

that was for my derivative part ..//

OpenStudy (anonymous):

that was my deravitive part that i squared and simplified...